Yurman Inc. uses a job-order costing system. During the month of May, the following transactions occurred:

May 1 Purchased materials on account for $29,740.
3 Requisitioned materials totaling $24,460 for use in production. Of the total, $9,190 was for Job 58, $8,910 for Job 59, and the remainder for Job 60.
31 Incurred direct labor for the month of $34,010, with an average wage of $19 per hour. Job 58 used 780 hours; Job 59, 610 hours; and Job 60, 400 hours.
31 Incurred and paid actual overhead of $17,850 (credit Various Payables).
31 Charged overhead to production at the rate of $4.70 per direct labor hour.
31 Completed and transferred Jobs 58 and 59 to Finished Goods.
31 Sold Job 57 (see beginning balance of Finished Goods) and Job 58 to their respective clients on account for a price of cost plus 40%.

Beginning balances as of May 1 were:

Materials $2,350
Work in Process 0
Finished Goods (Job 57) 25,580
Required:
1. Prepare journal entries for transactions from May 1 through 31.
2. Prepare brief job-order cost sheets for Jobs 58, 59, and 60.
3. Calculate the ending balance of Raw Materials.
4. Calculate the ending balance of Work in Process.
5. Calculate the ending balance of Finished Goods.

Chart of Accounts

CHART OF ACCOUNTS
Yurman Inc.
General Ledger
ASSETS
110 Cash
120 Accounts Receivable
131 Supplies
132 Prepaid Insurance
141 Raw Materials
142 Work in Process
143 Overhead Control
144 Finished Goods
170 Land
180 Equipment
181 Accumulated Depreciation
LIABILITIES
210 Accounts Payable
220 Wages Payable
230 Various Payables
EQUITY
310 Common Stock
320 Retained Earnings
REVENUE
410 Sales Revenue
EXPENSES
510 Cost of Goods Sold
520 Insurance Expense
530 Wages Expense
540 Supplies Expense
550 Utilities Expense
560 Depreciation Expense
590 Miscellaneous Expense

Amount Descriptions

Refer to the list below for the exact wording of text items within your job-order cost sheets.

Amount Descriptions
Applied overhead
Direct labor
Direct materials
Total

Journal

1. Prepare journal entries for transactions from May 1 through 31. Refer to the Chart of Accounts for the exact wording of account titles. Enter all amounts rounded to the nearest whole dollar.

PAGE 1

GENERAL JOURNAL

DATE ACCOUNT POST. REF. DEBIT CREDIT

1

2

3

4

5

6

7

8

9

10

11

12

13

14

15

16

Job-Order Cost Sheets

2. Prepare job-order cost sheets for Jobs 58, 59, and 60. Refer to the list of Amount Descriptions for the exact wording of text items within your job-order cost sheets. All amount cells must have an entry; enter zeros as needed.

Yurman Inc.

Job-Order Cost Sheets

For the Month of May

1

Job 58

Job 59

Job 60

2

3

4

5

Ending Balances

3. Calculate the ending balance of Raw Materials.

Raw Materials ending balance is .

4. Calculate the ending balance of Work in Process.

Work in Process ending balance is .

5. Calculate the ending balance of Finished Goods.

Finished Goods ending balance is .
Answer #1
GENERAL JOURNAL
DATE ACCOUNT POST. REF. DEBIT CREDIT
1. May 1 Raw materials $29740
2. Accounts payable $29740
3. May 3 Work in process $24460
4. Raw materials $24460
5. May 31 Work in process $34010
6. Wages payable $34010
7. May 31 Overhead control $17850
8. Various payable $17850
9. May 31 Work in process (780+610+400)*$4.70 $8413
10. Overhead control $8413
11. May 31 Finished goods $51043
12. Work in process $51043
13. May 31 Accounts receivable ($53256*1.40) $74558
14. Sales revenue $74558
15. Cost of goods sold $53256
16. Finished goods $53256

Finished goods= Job 58+Job 59= $27676+23367= $51043

Cost of goods sold= Job 57+Job 58= $25580+27676= $53256

Job 57= Beginning balance= $25580

Yurman Inc.
Job-Order Cost Sheets
For the Month of May
1. Amount Descriptions Job 58 Job 59 Job 60
2. Direct materials $9190 $8910 $6360
3. Direct labor (780*$19)= 14820 (610*$19)= 11590 (400*$19)= 7600
4. Applied overhead (780*$4.70)= 3666 (610*$4.70)= 2867 (400*$4.70)= 1880
5. Total $27676 $23367 $15840

Direct materials of Job 60= $24460-9190-8910= $6360

3) Raw materials ending balance= Beginning balance+Purchases-Used in production

= $2350+29740-24460= $7630

4) Work in process ending balance= Beginning balance+Direct materials+Direct labor+Applied overhead-Finished goods

= $0+24460+34010+8413-51043= $15840

5) Finished goods ending balance= Beginning balance+Cost of goods manufactured-Cost of goods sold

= $25580+51043-53256= $23367